solve the equation and round it to the nearest 3 decimal places if necessary.
ln(x-2)+3=6

Answer by ankor@dixie-net.com(22740)

ln(x-2)+3 = 6
ln(x-2) = 6 - 3
ln(x-2) = 3
Use exponent equiv
x - 2 = e^3
x - 2 = 20.08553
x = 20.08553 + 2
x = 22.08553 ~ 22.086